# Unified Text-Image-to-Video Generation: A Training-Free Approach to Flexible Visual Conditioning

### v2

@@ 5.1 Dataset and Metrics
Dataset We run experiments on the widely-used UCF-101 dataset [soomro2012ucf101] following [ho2022video, fu2023tell, ni2024ti2v], which is composed of 13,320 videos of 101 action categories. We select 42 action classes including complex actions with high dynamics (e.g., surfing) and small motion with still background (e.g., typing). For each class, we randomly select 25 instances (totaling 1,050 videos) as the test set. We manually rewrite the textual prompts for each action. Please also refer to Supp. Sec. B.2 for experiments on another synthetic dataset.
@@
Implementation Details For a fair comparison with [ni2024ti2v], we implement our FlexTI2V method on ModelScopeT2V [wang2023modelscope], which shows great robutness to out-domain T2V generation. We set the number of video frames as $M=16$, and total denoising steps $T=20$ with deterministic DDIM sampling (i.e., $\sigma_{t}=0$ in Eq. 2). The initial percentage and steps of swapping are set as $P_{0}=0.3$ and $t_{0}=10$. The decreasing strides are $\delta_{1}=5\times 10^{-3}$ and $\delta_{2}=0.3$. All experiments are conducted on 1 NVIDIA H100 GPU. Please refer to Supp. Sec. C.2 for more implementation details.
@@
Figure 5: Generated videos of our method implemented to Wan2.1-T2V. The images with blue edges are condition images and images with green edges are generated video frames. Please refer to the supplementary for corresponding videos and extra demos.
5.2 Quantitative Comparison
We compare with prior approaches quantitatively under three classic settings: (1) image animation – one condition image as the first video frame ($N=0$, $\bm{p}=\{0\}$), (2) rewinding – one condition image as the last video frame ($N=0$, $\bm{p}=\{15\}$), and (3) video frame interpolation – two condition images as the first and last frames ($N=2$, $\bm{p}=\{0,15\}$). We also introduce a new setting, video frame outpainting – one condition image as the middle video frame ($N=1$, $\bm{p}=\{7\}$). We mainly compare with two state-of-the-art training-free methods – TI2V-Zero [ni2024ti2v] and time reversal fusion (TRF) [feng2024explorative]. TI2V-Zero is initially designed only for image animation task. We adapt it to outpainting and rewinding tasks by adjusting the order of frame synthesis. Moreover, we also include DynamiCrafter [xing2024dynamicrafter] as a training-based baseline for a thorough comparison. More details of baseline implementation are elaborated in Supp. Sec. C.3. The results are shown in Tab. 1.
@@ 5.2 Quantitative Comparison
User Study Besides automatic metrics. We also conduct user study for a thorough evaluation. We sample 100 instances from our test set for each task, with each instance evaluated by five raters. Each rater is asked to select the video with best alignment with the textual prompt and condition images. The human preference percentage of each method is presented in Tab. 1. FlexTI2V is selected as the best in the highest percentage of samples compared with other competitors. The results suggest that the videos synthesized by our method aligns better with human’s subjective perception.
5.3 Visualization for Qualitative Comparison
Besides quantitative evaluation, we also visualize the videos synthesized by our method and other baselines for a qualitative comparison. As illustrated in Fig. 4, we observe some distortion in the videos generated by DynamiCrafter (e.g., the person in image animation and the knife in video frame interpolation), which may be the explanation for the low performance in Tab. 1. TRF is good at preserving visual features of the two bounded images. However, the video frames synthesized by TRF are blurry around the knife. We speculate that TRF is initially proposed for smooth motion generation and camera viewpoint transition. It may be weak in modeling object state change, such as the onion being cut into small pieces in this example. TI2V-Zero can generate videos with large motion because of the autoregressive generation mechanism. However, the cumulative mistakes weaken the controllability of visual conditions, leading to a low consistency in the frames far from the condition image. In contrast, our method can follow the textual prompts and visual conditions faithfully, and synthesize smooth motions in all TI2V settings. Please refer to Secs. B.6 and B.7 in the supplementary for more visualization of our method.
5.4 Generalization to Different Foundation Models
We use ModelScopeT2V [wang2023modelscope] as the foundation model in previous experiments simply for a fair comparison with prior methods. We also implement our method to another SOTA T2V foundation model – Wan2.1-T2V [wang2025wan] using transformer-based architectures. As shown in Fig. 5, our method can also incorporate additional image conditioning to Wan2.1 and synthesize videos with high fidelity and smooth motions, showing strong generalization capability to different T2V models. The quantitative comparison with Wan2.1 model family is demonstrated in Supp. Sec. B.1.
@@ 5.5 Ablation Study
In our method, there are two components (frame replacement and random patch swapping) to incorporate visual conditioning into T2V diffusion models. We evaluate the contribution of each component by ablating them from our method separately. The results are presented in Tab. 3. After removing frame replacement, the performance lags behind the full model by a small margin in all tasks. The result indicates that random patch swapping (RPS) can work alone to effectively merge visual guidance into T2V diffusion process. It also validates our hypothesis that sufficient visual features can be transferred through a portion of local patches at each denoising step for video generation. On the contrary, if we remove RPS and only implement frame replacement strategy for extra image conditioning, a significant performance drop is observed across all metrics and tasks. Though frame replacement introduces image features into the temporal layers, the model estimates the holistic noise added to all video frames (16 frames in our experimental setting) as an entirety. Hence, the condition image features are likely to be ignored because of the small number of images against the large number of video frames. We observe that the videos generated without RPS greatly deviate from the condition images, thus resulting in a remarkable decrease in the performance. More ablation and hyperparameter analysis are provided in Supp. Secs. B.3, B.4 and B.5.
@@ 6 Conclusion
In this paper, we introduce a generic text-image-to-video (TI2V) problem that unifies existing TI2V tasks, which requires the model be able to synthesize smooth videos conditioned on an arbitrary number of images at any expected positions. To solve this problem, we propose a novel training-free method, FlexTI2V, that can introduce flexible image conditioning to pretrained text-to-video diffusion models. The key component of our method is random patch swapping strategy, which merges visual features of condition images into video denoising process by swapping a random portion of local patches. The extensive experiments indicate a notable improvement of FlexTI2V over prior methods, as well as the flexibility of our model. Our work is an important step for flexible and low-cost controllable video generation, providing an alternative way to adapt pretrained foundation models to specific downstream tasks.
[…]
@@
This is the supplementary material for the publication titled “Unified Text-Image-to-Video Generation: A Training-Free Approach to Flexible Visual Conditioning”. We organize the content as follows:
@@
B.1 – Quantitative Comparison with Wan2.1 Model Family
B.2 – Quantitative Results on Synthetic Dataset
B.3 – Analysis of Dynamic Control Strategy
B.4 – Visualization for Ablation Study
B.5 – Ablation of Hyperparameters
B.6 – Generated Videos without Humans
B.7 – Additional Demonstration of Our Method
B.8 – Failure Cases
@@
8:  $\hat{\bm{z}}_{t}\leftarrow\verb+FrameReplace+(\tilde{\bm{x}}_{t},\ \bm{z}_{t},\ \bm{p})$ $\triangleright$ Frame Replacement, Eq.(3) in Main Paper
@@
A Algorithm Formulation
We formulate our FlexTI2V approach in Alg. A. The function $\verb+Repeat+(\tilde{x}_{T}^{(0)},M)$ at step 3 duplicates the inverted condition image representation for $M$ times as the initialization for video frame denoising. If there are more than one condition images, we find using different image makes trivial difference in the generation performance. We use the first condition image in our method. The function $\verb+BoundIndex+(m,\bm{p})$ at step 8 returns the indices of the two closest condition images before and after $m$-th video frame respectively (i.e., bounded index). If all condition images are before or after this video frame, the function returns only one index of the closest image.
B Additional Experiment Resutls
B.1 Quantitative Comparison with Wan2.1 Model Family
We additionally provide quantitative comparison with Wan2.1-I2V in image animation and Wan2.1-FLF2V in video frame interpolation on UCF-101 [soomro2012ucf101]. As shown in Tab. A and Tab. B Our method can be successfully applied to the SOTA T2V models and achieve a better performance compared with the finetuned counterpart. The possible reason is that it’s still challenging for training-based methods to generalize to out-of-distribution data (e.g., different aspect ratios and resolutions). This also shows the necessity of training-free methods like our work.
[…]
Wan2.1-I2V [wang2025wan] 141.22 31.01 82.33
Wan2.1-T2V + Our Method 101.74 30.98 85.87
Table A: Comparison with Wan2.1-I2V in image animation. The orange column refers to our method.
[…]
Wan2.1-FLF2V [github2025wan] 112.22 31.47 85.13
Wan2.1-T2V + Our Method 90.82 31.50 86.04
Table B: Comparison with Wan2.1-FLF2V in video interpolation. The orange column refers to our method.
@@ B Additional Experiment Resutls
Table C: Results on the synthetic dataset. The orange column refers to our method. Refer to Sec. B.2 for more descriptions.
B.2 Quantitative Results on Synthetic Dataset
@@
As shown in Tab. C, our method still prominently outperforms all the baseline approaches in CLIP-T (text-video alignment) and CLIP-I (image-video similarity). The results validate that our method is also superior in modeling different types of motions and videos in the TI2V task.
Figure A: Demonstration of ablation study. The images with blue edges are condition images and images with green edges are generated video frames. More explanations are provided in Sec. B.3 and Sec. B.4.
@@
Figure B: Additional examples of our method on videos without humans. The images with blue edges are condition images and images with green edges are generated video frames. Please see Sec. B.6 for more details.
B.3 Analysis of Dynamic Control Strategy
We remove the dynamic control strategy from our method and use a fixed percentage and steps for random patch swapping. The results are illustrated in the last row of Fig. A. The motion in the generated video frames is trivial without dynamic control. The reason is because the constraints of visual conditioning are overly strong for all frames. The result validates the necessity of the dynamic control strategy.
B.4 Visualization for Ablation Study
Besides ablation of dynamic control (see Sec. B.3), we also show additional visualization for quantitative ablation study of patch replacement and random patch swapping in Fig. A. After removing patch replacement, the output video is still decent with minor deviation from the condition image. However, without random patch swapping, the generated video fails to follow the visual condition, thus leading to a poor performance in main paper Tab. 3.
B.5 Analysis of Hyperparameters
In our experiments, the optimal hyperparameters are selected empirically. Below are the ablation studies for $\delta_{1}$, $\delta_{2}$, $P_{0}$ and $t_{0}$ on the image animation task. We change only one hyperparameter in each table.
As shown in Tab. D, if we use a large $\delta_{1}$ and $\delta_{2}$, or small $P_{0}$ and $t_{0}$, the image conditioning is weaker than the optimal hyperparameters so that the generated video can not accurately follow the images. Likewise, if $\delta_{1}$ and $\delta_{2}$ are small, or $P_{0}$ and $t_{0}$ are large, the image conditioning is too strong so that the output videos are nearly frozen. To achieve an optimal balance, we select the best hyperparameter combination based on the ablation study.
$\delta_{1}$ $4\times 10^{-3}$ $\bm{5\times 10^{-3}}$ $6\times 10^{-3}$ $7\times 10^{-3}$
[…]
$\delta_{2}$ $0.2$ $\bm{0.3}$ $0.4$ $0.5$
[…]
   $P_{0}$    $0.2$    $\bm{0.3}$    $0.4$
[…]
   $t_{0}$    77    $10$    $13$
[…]
Table D: Hyperparameter analysis. Please refer to Sec. B.5 for more details.
B.6 Generated Videos without Humans
The videos in UCF-101 are mainly about human actions. To assess the generalization performance of our approach, we illustrate the demonstration of FlexTI2V synthesizing videos without humans in main paper Fig. 5. We supplement extra examples in Fig. B. Our approach seamlessly generalizes to the motion of objects showing strong generalization capability.
Figure C: Additional demonstration of videos synthesized by FlexTI2V in classic TI2V settings. The images with blue and pink edges are condition imags, and images with green edges are generated video frames. NN and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
Figure D: Additional demonstration of videos synthesized by FlexTI2V in new TI2V settings. The images with blue, pink and orange edges are condition imags, and images with green edges are generated video frames. NN and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
B.7 Additional Demonstration of Our Method
We illustrate more videos synthesized by our method in classic TI2V generation settings (see Fig. C) and new settings (see Fig. D). Our approach is able to condition the pretrained T2V diffusion foundation model on any number of images with flexible positions.
@@ B.7 Additional Demonstration of Our Method
Figure E: Typical failure cases of our approach. The images with blue and pink edges are condition images and images with green edges are generated video frames. Please refer to Sec. B.8 for more analysis.
B.8 Failure Cases

### v3

@@ 5.1 Dataset and Metrics
Dataset We run experiments on the widely-used UCF-101 dataset [soomro2012ucf101] following [ho2022video, fu2023tell, ni2024ti2v], which is composed of 13,320 videos of 101 action categories. We select 42 action classes including complex actions with high dynamics (e.g., surfing) and small motion with still background (e.g., typing). For each class, we randomly select 25 instances (totaling 1,050 videos) as the test set. We manually rewrite the textual prompts for each action. We also run experiments on one more synthetic dataset to consolidate our conclusion, which is shown in the supplementary.
@@
Implementation Details For a fair comparison with [ni2024ti2v], we implement our FlexTI2V method on ModelScopeT2V [wang2023modelscope], which shows great robutness to out-domain T2V generation. We set the number of video frames as $M=16$, and total denoising steps $T=20$ with deterministic DDIM sampling (i.e., $\sigma_{t}=0$ in Eq. 2). The initial percentage and steps of swapping are set as $P_{0}=0.3$ and $t_{0}=10$. The decreasing strides are $\delta_{1}=5\times 10^{-3}$ and $\delta_{2}=0.3$. Our method can be run on 1 NVIDIA H100 GPU. Please refer to the supplementary for more implementation details.
[…]
Image Animation
DynamiCrafter [xing2024dynamicrafter] 26.20 210.77 832.38 29.86 89.10 80.95
TI2V-Zero [ni2024ti2v] 28.40 189.49 586.54 29.84 88.01 82.29
FlexTI2V 45.40 125.49 558.29 30.97 94.40 84.38
Rewinding
TI2V-Zero [ni2024ti2v] 39.80 199.01 611.60 29.77 89.21 81.02
FlexTI2V 60.20 134.98 596.96 30.95 94.58 85.11
Outpainting
TI2V-Zero [ni2024ti2v] 38.60 216.80 641.07 28.80 89.01 85.01
FlexTI2V 61.40 167.46 632.71 30.75 95.89 88.34
Video Frame Interpolation
DynamiCrafter [xing2024dynamicrafter] 16.20 310.74 731.90 29.69 84.26 78.68
TRF [feng2024explorative] 32.20 247.88 631.54 29.47 91.99 83.35
FlexTI2V 51.60 98.41 406.30 30.92 92.34 86.88
Table 1: Quantitative comparison with previous methods. “User” denotes the user study results. $\downarrow$ means a lower score indicates better performance. The orange row refers to our method.
5.2 Quantitative Comparison
We compare with prior approaches quantitatively under three classic settings: (1) image animation – one condition image as the first video frame ($N=0$, $\bm{p}=\{0\}$), (2) rewinding – one condition image as the last video frame ($N=0$, $\bm{p}=\{15\}$), and (3) video frame interpolation – two condition images as the first and last frames ($N=2$, $\bm{p}=\{0,15\}$). We also introduce a new setting, video frame outpainting – one condition image as the middle video frame ($N=1$, $\bm{p}=\{7\}$). We mainly compare with two state-of-the-art training-free methods – TI2V-Zero [ni2024ti2v] and time reversal fusion (TRF) [feng2024explorative]. TI2V-Zero is initially designed only for image animation task. We adapt it to outpainting and rewinding tasks by adjusting the order of frame synthesis. Moreover, we also include DynamiCrafter [xing2024dynamicrafter] as a training-based baseline for a thorough comparison. More details of baseline implementation are elaborated in the supplementary. The results are shown in Tab. 1.
DynamiCrafter [xing2024dynamicrafter] TI2V-Zero [ni2024ti2v] TRF [feng2024explorative] FlexTI2V
Inference Time (s) 9.45 275.14 60.05 2.71
Table 2: Comparison of model efficiency in inference. The orange column refers to our method. Our method shows great advantage in inference speed compared with other baseline models.
User Study Besides automatic metrics. We also conduct user study for a thorough evaluation. We sample 100 instances from our test set for each task, with each instance evaluated by five raters. Each rater is asked to select the video with best alignment with the textual prompt and condition images. The human preference percentage of each method is presented in Tab. 1. FlexTI2V is selected as the best in the highest percentage of samples compared with other competitors. The results suggest that the videos synthesized by our method align better with human’s subjective perception.
@@ 5.2 Quantitative Comparison
5.3 Visualization for Qualitative Comparison
Besides quantitative evaluation, we also visualize the videos synthesized by our method and other baselines for a qualitative comparison. As illustrated in Fig. 4, we observe some distortion in the videos generated by DynamiCrafter (e.g., the person in image animation and the knife in video frame interpolation), which may be the explanation for the low performance in Tab. 1. TRF is good at preserving visual features of the two bounded images. However, the video frames synthesized by TRF are blurry around the knife. We speculate that TRF is initially proposed for smooth motion generation and camera viewpoint transition. It may be weak in modeling object state change, such as the onion being cut into small pieces in this example. TI2V-Zero can generate videos with large motion because of the autoregressive generation mechanism. However, the cumulative mistakes weaken the controllability of visual conditions, leading to a low consistency in the frames far from the condition image. In contrast, our method can follow the textual prompts and visual conditions faithfully, and synthesize smooth motions in all TI2V settings. More examples of our method are illustrated in the supplementary.
[…]
Image Animation
Wan2.1-I2V [wang2025wan] 141.22 31.01 82.33
Wan2.1-T2V + Our Method 101.74 30.98 85.87
Video Frame Interpolation
Wan2.1-FLF2V [github2025wan] 112.22 31.47 85.13
Wan2.1-T2V + Our Method 90.82 31.50 86.04
Table 3: Comparison with Wan2.1 model family. $\downarrow$ means a lower score in this metric indicates a better performance. The orange column refers to our method.
Figure 5: Generated videos of our method implemented to Wan2.1-T2V. The images with blue edges are condition images and images with green edges are generated video frames. Please refer to supplementary for corresponding videos and extra demos.
5.4 Generalization to Different Foundation Models
We use ModelScopeT2V [wang2023modelscope] as the foundation model in previous experiments simply for a fair comparison with prior methods. To validate the generalization of our method, we also implement our method to another SOTA T2V foundation model – Wan2.1-T2V [wang2025wan] using transformer-based architectures. We directly compare with Wan2.1-I2V and Wan2.1-FLF2V on image animation and video frame interpolation separately. The two specialized models are finetuned from the base Wan2.1-T2V model for the two specific tasks, which makes it an apple-to-apple comparison with our proposed approach.
As shown in Tab.˜3, our method outperforms the two specialized models on the two tasks. The possible reason is that it’s still challenging for training-based methods to generalize to out-of-distribution data (e.g., different aspect ratios and resolutions). This also shows the necessity of training-free methods like our work. We demonstrate videos synthesized by our method in Fig.˜5. Implemented on top of the modern architecture, our method generates videos with high fidelity and smooth motions, showing robust generalization to different T2V models.
@@ 5.5 Ablation Study
In our method, there are two components (frame replacement and random patch swapping) to incorporate visual conditioning into T2V diffusion models. We evaluate the contribution of each component by ablating them from our method separately. The results are presented in Tab. 4. After removing frame replacement, the performance lags behind the full model by a small margin in all tasks. The result indicates that random patch swapping (RPS) can work alone to effectively merge visual guidance into T2V diffusion process. It also validates our hypothesis that sufficient visual features can be transferred through a portion of local patches at each denoising step for video generation. On the contrary, if we remove RPS and only implement frame replacement strategy for extra image conditioning, a significant performance drop is observed across all metrics and tasks. Though frame replacement introduces image features into the temporal layers, the model estimates the holistic noise added to all video frames (16 frames in our experimental setting) as an entirety. Hence, the condition image features are likely to be ignored because of the small number of images against the large number of video frames. We observe that the videos generated without RPS greatly deviate from the condition images, thus resulting in a remarkable decrease in the performance. We also visualize examples of the ablation study for qualitative analysis. Please check out our supplementary.
$P_{0}$ $0.2$ $\bm{0.3}$ $0.4$
[…]
$\delta_{1}$ $4\times 10^{-3}$ $\bm{5\times 10^{-3}}$ $6\times 10^{-3}$ $7\times 10^{-3}$
[…]
$t_{0}$ 77 $\bm{10}$ $13$
[…]
$\delta_{2}$ $0.2$ $\bm{0.3}$ $0.4$ $0.5$
[…]
Table 5: Hyperparameter analysis. We change only one hyperparameter in each subtable and keep the other hyperparameters fixed. $\downarrow$ means a lower score indicates better performance. The orange columns refer to the hyperparameters we finally adopt.
5.6 Hyperparameter Analysis
In the proposed RPS approach, there are four important hyperparameters – initial swapping percentage $P_{0}$, initial number of swapping steps $t_{0}$ and the decreasing strides $\delta_{1}$, $\delta_{2}$ for swapping percentage and steps. The optimal hyperparameters are selected empirically based on the analysis in Tab.˜5. If we use a large $\delta_{1}$ and $\delta_{2}$, or small $P_{0}$ and $t_{0}$, the image conditioning is weaker than the optimal hyperparameters so that the generated video can not accurately follow the images. Likewise, if $\delta_{1}$ and $\delta_{2}$ are small, or $P_{0}$ and $t_{0}$ are large, the image conditioning is too strong so that the output videos are nearly frozen. To achieve an optimal balance, we select the best hyperparameter combination based on the experiments. We also find the optimal combination is very consistent across different datasets, so we use it in all the experiments.
@@ 6 Conclusion
In this paper, we introduce a generic text-image-to-video (TI2V) problem that unifies existing TI2V tasks, which requires the model be able to synthesize smooth videos conditioned on an arbitrary number of images at any expected positions. To solve this problem, we propose a novel training-free method, FlexTI2V, that can introduce flexible image conditioning to pretrained text-to-video diffusion models. The key component of our method is random patch swapping strategy, which merges visual features of condition images into video denoising process by swapping a random portion of local patches. The extensive experiments indicate a notable improvement of FlexTI2V over prior methods, as well as the flexibility of our model. Our work is an important step for flexible and low-cost controllable video generation, which provides an alternative way to adapt pretrained foundation models to specific downstream tasks, thus democratizing the access to customized video generation tools.
A Training-Free Approach to Flexible Visual Conditioning
@@
This is the supplementary material for the paper titled “Unified Text-Image-to-Video Generation: A Training-Free Approach to Flexible Visual Conditioning”. We organize the content as follows:
@@
B.1 – Quantitative Results on Synthetic Dataset
B.2 – Analysis of Dynamic Control Strategy
B.3 – Visualization for Ablation Study
B.4 – Generated Videos without Humans
B.5 – Additional Demonstration of Our Method
B.6 – Failure Cases
@@
8:  $\hat{\bm{z}}_{t}\leftarrow\verb+FrameReplace+(\tilde{\bm{x}}_{t},\ \bm{z}_{t},\ \bm{p})$ $\triangleright$ Frame Replacement, Eq. 3 in Main Paper
@@
Appendix A Algorithm Formulation
We formulate our FlexTI2V approach in Alg. A. The function $\texttt{Repeat}(\tilde{x}_{T}^{(0)},M)$ at step 3 duplicates the inverted condition image representation for $M$ times as the initialization for video frame denoising. If there are more than one condition images, we find using different image makes trivial difference in the generation performance. We use the first condition image in our method. The function $\verb+BoundIndex+(m,\bm{p})$ at step 8 returns the indices of the two closest condition images before and after $m$-th video frame respectively (i.e., bounded index). If all condition images are before or after this video frame, the function returns only one index of the closest image.
@@ Appendix A Algorithm Formulation
Table A: Results on the synthetic dataset. The orange column refers to our method. Refer to Sec.˜B.1 for more descriptions.
Appendix B Additional Experiment Resutls
B.1 Quantitative Results on Synthetic Dataset
@@
As shown in Tab.˜A, our method still prominently outperforms all the baseline approaches in CLIP-T (text-video alignment) and CLIP-I (image-video similarity). The results validate that our method is also superior in modeling different types of motions and videos in the TI2V task.
@@
Figure A: Additional examples of our method on videos without humans. The images with blue edges are condition images and images with green edges are generated video frames. Please see Sec. B.4 for more details.
@@
Figure B: Demonstration of ablation study. The images with blue edges are condition images and images with green edges are generated video frames. More explanations are provided in Sec. B.2 and Sec. B.3.
B.2 Analysis of Dynamic Control Strategy
We remove the dynamic control strategy from our method and use a fixed percentage and steps for random patch swapping. The results are illustrated in the last row of Fig. B. The motion in the generated video frames is trivial without dynamic control. The reason is because the constraints of visual conditioning are overly strong for all frames. The result validates the necessity of the dynamic control strategy.
B.3 Visualization for Ablation Study
Besides ablation of dynamic control (see Sec. B.2), we also show additional visualization for quantitative ablation study of patch replacement and random patch swapping in Fig. B. After removing patch replacement, the output video is still decent with minor deviation from the condition image. However, without random patch swapping, the generated video fails to follow the visual condition, thus leading to a poor performance in main paper Tab. 4.
Figure C: Typical failure cases of our approach. The images with blue and pink edges are condition images and images with green edges are generated video frames. Please refer to Sec. B.6 for more analysis.
B.4 Generated Videos without Humans
The videos in UCF-101 are mainly about human actions. To assess the generalization performance of our approach, we supplement extra examples of FlexTI2V synthesizing videos without humans in Fig. A. Our approach seamlessly generalizes to the motion of objects showing strong generalization capability.
Figure D: Additional demonstration of videos synthesized by FlexTI2V in classic TI2V settings. The images with blue and pink edges are condition imags, and images with green edges are generated video frames. NN and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
Figure E: Additional demonstration of videos synthesized by FlexTI2V in new TI2V settings. The images with blue, pink and orange edges are condition imags, and images with green edges are generated video frames. NN and $\bm{p}$ are the number and positions of condition images. Our method generates videos with 16 frames in the experiments.
B.5 Additional Demonstration of Our Method
We illustrate more videos synthesized by our method in classic TI2V generation settings (see Fig. D) and new settings (see Fig. E). Our approach is able to condition the pretrained T2V diffusion foundation model on any number of images with flexible positions.
B.6 Failure Cases